# Discovering Effective Policies for Land-Use Planning

### 13 Dec 2023 v2

@@
Abstract
How areas of land are allocated for different uses, such as forests, urban, and agriculture, has a large effect on carbon balance, and therefore climate change. Based on available historical data on changes in land use and a simulation of carbon emissions/absorption, a surrogate model can be learned that makes it possible to evaluate the different options available to decision-makers efficiently. An evolutionary search process can then be used to discover effective land-use policies for specific locations. Such a system was built on the Project Resilience platform (ITU, 2023) and evaluated with the Land-Use Harmonization dataset (Hurtt et al., 2020) and the BLUE simulator (Hansis et al., 2015). It generates Pareto fronts that trade off carbon impact and amount of change customized to different locations, thus providing a potentially useful tool for land-use planning.
1 Introduction
One of the main factors contributing to climate change is how much land area is allocated for different uses (Friedlingstein et al., 2022). Forests in general absorb more carbon than e.g. crops and ranges, yet such uses are essential for the economy. Land-use patterns must therefore be planned to maximize carbon absorption while maintaining economic viability.
@@ 1 Introduction
More specifically, the tool is designed to answer three questions: (1) For a geographical grid cell, identified by its latitude and longitude, what changes to the land usage can be made to reduce CO2 emissions? (2) What will be the long-term CO2 impact of changing land usage in a particular way? (3) What are the optimal land-use choices that can be made with minimal cost and maximal effect?
The approach is based on the Evolutionary Surrogate-assisted Prescription method (Francon et al., 2020). The idea is to first utilize historical data to learn a surrogate model on how land-use decisions in different contexts affect carbon emissions. Then, this model is used to evaluate candidates in an evolutionary search process for good land-use change policies.
@@
An interactive demo of the system is at https://landuse.evolution.ml.
2 Background
2.1 Evolutionary Surrogate-assisted Prescription
Figure 1: The ESP Decision Optimization Method. A predictor is trained with historical data on how given actions in given contexts led to specific outcomes. It is then used as a surrogate in order to evolve prescriptors, i.e. neural networks that implement decision policies resulting in the best possible outcomes.
@@ 2.1 Evolutionary Surrogate-assisted Prescription
such that $\sum_{i,j}O_{j}(C_{i},A_{i})$ over all possible contexts $i$ and outcome dimensions $j$ is maximized (assuming they improve with increase). It thus approximates the optimal decision policy for the problem. The predictor can be learned from historical data on $CAO$ triples. In contrast, the optimal actions $A$ for each context $C$ are not known, and must therefore be found through search.
@@
ESP was first evaluated in reinforcement learning tasks such as function approximation, cart-pole control, and the flappybird game, and found to discover significantly better solutions, find them faster, and with lower regret than standard approaches such as direct evolution, DQN, and PPO. Its power comes from automatic regularization and automatic curricular learning: When predictors are learned at the same time as the prescriptors, each prescriptor is evaluated against multiple predictors; as the predictors improve, they provide more refined evaluations. ESP is thus a powerful mechanism for learning policies in time-varying domains.
2.2 Project Resilience
@@
A major application of ESP to decision-making was developed for optimizing strategies for non-pharmaceutical interventions (NPIs) in the COVID-19 pandemic (Miikkulainen et al., 2021). Using case and NPI data from around the world, collected in 2020-2022 by Oxford University’s Covid-19 project (Hale et al., 2020), a neural network model was developed to predict the course of the pandemic; this network was then used as a surrogate in evolving another neural network to prescribe NPIs. It discovered different strategies at different phases of the pandemic, such as focusing on schools and workplaces early on, alternating policies over time, and focusing on public information and masking later. This application demonstrated the power of ESP on discovering good tradeoffs for complex real-world decision-making tasks.
@@
Encouraged by this application, the XPRIZE Pandemic Response Challenge (XPRIZE, 2023) was developed in 2020-2021 for NPI optimization using the same predictor+prescriptor approach. The competitors used several different machine learning or other techniques, and drew significant attention to AI-supported decision-making. The results were effective and in some cases deployed by decision-makers. An important conclusion was that it was possible to convene a group of experts – data scientists, epidemiologists, and public health officials, to build a useful set of tools to advise us on how to cope with and plan around a health disruption to society.
@@
This experience with the XPRIZE Pandemic Response Challenge then led to Project Resilience (ITU, 2023). The XPRIZE competition serves as a blueprint for a collaborative effort to tackle global problems. The goal of Project Resilience is to build a public AI utility where a global community of innovators and thought leaders can enhance and utilize a collection of data and AI approaches to help with better preparedness, intervention, and response to environmental, health, information, social equity, and similarly scoped problems. The project is hosted by the ITU agency of the United Nations, under its Global Initiative for AI and Data Commons, and contributes to its general efforts towards meeting the Sustainable Development Goals (SDGs).
@@
  • Build an API for third parties to submit models.
@@
  • Build a portal to visualize assessment of predictors and prescriptors to include generations of key performance indicators (KPIs) and comparison across models;
@@
3 Land-Use Optimization Task
@@
3.1 Data
The data for carbon emissions (Emissions resulting from Land-Use Change, ELUC) originate from a high-fidelity simulation called Bookkeeping of Land-Use Emissions (BLUE; Hansis et al., 2015). BLUE is designed to estimate the long-term CO2 impact of committed land use. “Committed emissions” means all the emissions that are caused by a land-use change event are attributed to the year of the event. BLUE is a bookkeeping model that attributes carbon fluxes to land-use activities. While in principle a simulator can be used as the surrogate model for ESP, in practice the simulations are too expensive to carry out on demand during the search for good policies. Therefore, the team in charge of the BLUE model performed a number of simulations covering a comprehensive set of situations for 1850-2022, resulting in a dataset that could be used to train an efficient surrogate model.
The Land-Use Change (LUC) data is provided by the Land-Use Harmonization project (LUH2; Hurtt et al., 2020). A land-use harmonization strategy estimates the fractional land-use patterns, underlying land-use transitions, and key agricultural management information, annually for the time period 850-2100 at 0.25 x 0.25 degree resolution.
The most recent version of the land-use data (2022) consists of separate layers for states, transitions, and management. These files are based on the new HYDE3.3 historical database, as well as the 2021 FAO wood harvest data, for all years 850-2022. It consists of the following land-use types:
@@ 3.1 Data
  • c3ann: Annual C3 crops (e.g. wheat)
  • c4ann: Annual C4 crops (e.g. maize)
  • c3per: Perennial C3 crops (e.g. banana)
  • c4per: Perennial C4 crops (e.g. sugarcane)
  • c3nfx: Nitrogen fixing C3 crops (e.g. soybean)
@@
  • range: Natural grassland / savannah / desert / etc.
@@
3.2 Decision-making Problem
The modeling approach aims to understand the domain in two ways: (1) In a particular situation, what are the outcomes of the decision maker’s actions? (2) What are the actions that result in the best outcomes, i.e. the lowest carbon emission (or highest capture) with the smallest change? The data is thus organized into context, action, and outcome variables.
Context describes the problem the decision maker is facing, i.e. a particular grid cell, a point in time when the decision has to be made, and the usage of the land at that point. More specifically it consists of:
@@ 3.2 Decision-making Problem
  • Land usage, representing the percentage of the land used by each land type above. In addition to those listed above, there is a ‘nonland’ type that represents the percentage of the cell that is not land (e.g. typically sea, lake, etc.). The land usage types sum up to 100%.
@@
First, decision-makers cannot affect primary land. The idea is that it is always better to preserve primary vegetation; destroying it is not an option given to the system. Technically, it is not possible to re-plant primary vegetation. Once destroyed, it is destroyed forever. If re-planted, it would become secondary vegetation.
@@
Second, decision-makers cannot affect urban areas. The needs of urban areas are dictated by other imperatives, and optimized by other decision makers. Therefore, the system cannot recommend that a city should be destroyed, or expanded.
@@
Outcomes consist of two conflicting variables. The primary variable is ELUC, i.e. emissions from land-use change. It consists of all CO2 emissions attributed to the change, in metric tons of carbon per hectare (tC/ha), obtained from the BLUE simulation. A positive number means carbon is emitted, a negative number means carbon is captured. The secondary variable is the cost of the change, represented by the percentage of land that was changed. This variable is calculated directly from the actions.
@@
There is a trade-off between these two objectives: It is easy to reduce emissions by changing most of the land, but that would come at a huge cost. Therefore, decision-makers have to minimize ELUC while minimizing land change at the same time. Consequently, the result is not a single recommendation, but a Pareto front where each point represents the best implementation of each tradeoff given a balance between the two outcomes.
4 Models
@@
The system consists of the predictor, trained with supervised learning on the historical data, and the prescriptor, trained through evolution.
Prediction:
@@
Given the context and actions that were performed, the predictive model estimates the outcomes. In this case, since the cost outcome can be calculated directly, only the ELUC is predicted by the model. That is, given the land usage of a specific location, and the changes that were made during a specific year, the model predicts the CO2 long-term emissions directly caused by these changes.
@@
Any predictive model can be used in this task, including a neural network, random forest, or linear regression. As usual, the model is fit to the existing historical data and evaluated with left-out data.
Prescription:
@@
Given context, the prescriptive model suggests actions that optimize the outcomes. The model has to do this for all possible contexts, and therefore it represents an entire strategy for optimal land use. The strategy can be implemented in various ways, including decision trees, sets of rules, or neural networks. The approach in this paper is based on neural networks.
@@ 5 Experiments
The main (early) results in the project so far, on different ways to create good predictors, on evolving effective prescriptors, and on an interactive interface for analyzing the results, are described in this section, with the next steps outlined in Section 6.
5.1 Prediction
Model Train Time(s) W. Europe South America United States Global
LinReg (EU) 0.1236 0.0331 0.2570 0.1720 0.2204
LinReg (SA) 0.8281 0.1422 0.1549 0.0648 0.1110
LinReg (US) 0.5652 0.1402 0.1467 0.0345 0.0720
LinReg (Global) 11.1449 0.1410 0.1520 0.0366 0.0723
RF (EU) 43.2344 0.0314 0.2373 0.1156 0.2232
RF (SA) 338.2790 0.1722 0.0715 0.0462 0.1095
RF (US) 115.4622 0.1559 0.1874 0.0200 0.1120
RF (Global) 255.3720 0.1007 0.0870 0.0257 0.0558
Neural Net (EU) 78.7409 0.0247 0.5510 0.3313 0.3493
Neural Net (SA) 1208.0376 0.7950 0.0936 0.1866 0.2195
Neural Net (US) 312.3967 0.3348 0.2164 0.0196 0.1319
Neural Net (Global) 186.2248 0.3298 0.2182 0.1309 0.0530
Table 1: Mean absolute error of the models trained on each region, evaluated on that region as well as all other regions. The Neural Net models are more accurate than the RF models (with the exception of SA), which are more accurate than LinReg, and the specialized models are more accurate than the global model. The Neural Net models both capture nonlinearities and extrapolate well in extreme cases; therefore, the Global Neural Net model was used to evolve the prescriptors.

### 09 Feb 2024 v3

@@
Abstract.
How areas of land are allocated for different uses, such as forests, urban areas, and agriculture, has a large effect on the terrestrial carbon balance, and therefore climate change. Based on available historical data on land-use changes and a simulation of the associated carbon emissions and removals, a surrogate model can be learned that makes it possible to evaluate the different options available to decision-makers efficiently. An evolutionary search process can then be used to discover effective land-use policies for specific locations. Such a system was built on the Project Resilience platform (ITU, 2023) and evaluated with the Land-Use Harmonization dataset LUH2 (Hurtt et al., 2020) and the bookkeeping model BLUE (Hansis et al., 2015). It generates Pareto fronts that trade off carbon impact and amount of land-use change customized to different locations, thus providing a potentially useful tool for land-use planning.
Key words and phrases:
Surrogate modeling, Decision making, Climate Change, Land-use Change, Neural Networks, Evolutionary Computation
{CCSXML}
<ccs2012> <concept> <concept_id>10010147.10010257.10010321</concept_id> <concept_desc>Computing methodologies Machine learning algorithms</concept_desc> <concept_significance>500</concept_significance> </concept> <concept> <concept_id>10010147.10010919.10010172</concept_id> <concept_desc>Computing methodologies Distributed algorithms</concept_desc> <concept_significance>100</concept_significance> </concept> <concept> <concept_id>10010147.10010257.10010293.10011809.10011812</concept_id> <concept_desc>Computing methodologies Genetic algorithms</concept_desc> <concept_significance>500</concept_significance> </concept> <concept> <concept_id>10010147.10010257.10010293.10010294</concept_id> <concept_desc>Computing methodologies Neural networks</concept_desc> <concept_significance>500</concept_significance> </concept> </ccs2012>
1. Introduction
One of the main factors contributing to climate change are CO2 emissions due to land-use change (Friedlingstein et al., 2023). Land-use CO2 emissions crucially depend on how much land area is allocated for different uses. Forests in general take up more carbon than e.g. croplands and rangelands, yet such uses are essential for the economy. Land-use patterns must therefore be planned to maximize carbon removals while maintaining economic viability.
@@ 1. Introduction
More specifically, the tool is designed to answer three questions: (1) For a geographical grid cell, identified by its latitude and longitude, what changes to the land use can be made to reduce CO2 emissions? (2) What will be the long-term CO2 impact of changing land use in a particular way? (3) What are the optimal land-use choices that can be made with minimal cost and maximal effect?
The approach is based on the Evolutionary Surrogate-assisted Prescription method (Francon et al., 2020). The idea is to first utilize historical data to learn a surrogate model on how land-use decisions in different contexts affect carbon emissions, and then use this model to evaluate candidates in an evolutionary search process for good land-use change policies.
@@
Project code and data are available at https://github.com/<anony-mous>, and an interactive demo at https://landuse.evolution.ml.
2. Background
The surrogate-assisted evolutionary optimization method used on the project is first outlined below, followed by the project background as the first application of Project Resilience.
2.1. Evolutionary Surrogate-assisted Prescription
Figure 1. The Evolutionary Surrogate-assisted Prescription (ESP) method for decision optimization. A predictor is trained with historical data on how given actions in given contexts led to specific outcomes. It is then used as a surrogate in order to evolve prescriptors, i.e. neural networks that implement decision policies resulting in the best possible outcomes.
@@ 2.1. Evolutionary Surrogate-assisted Prescription
such that $\sum_{i,j}O_{j}(C_{i},A_{i})$ over all possible contexts $i$ and outcome dimensions $j$ is maximized (assuming increased outcomes are better). It thus approximates the optimal decision policy for the problem. The predictor can be learned from historical data on $CAO$ triples. In contrast, the optimal actions $A$ for each context $C$ are not known, and must therefore be found through search.
@@
ESP was first evaluated in reinforcement learning tasks such as function approximation, cart-pole control, and the flappybird game, and found to discover significantly better solutions, find them faster, and with lower regret than standard approaches such as direct evolution, Deep Q-Networks (DQN; Mnih et al., 2015), and Proximal Policy Optimization (PPO; Schulman et al., 2017). Its power comes from automatic regularization and automatic curricular learning: When predictors are learned at the same time as the prescriptors, each prescriptor is evaluated against multiple predictors; as the predictors improve, they provide more refined evaluations. ESP is thus a powerful mechanism for learning policies in time-varying domains.
2.2. Project Resilience
@@
A major application of ESP to decision-making was developed for optimizing strategies for non-pharmaceutical interventions (NPIs) in the COVID-19 pandemic (Miikkulainen et al., 2021). Using case and NPI data from around the world, collected in 2020-2022 by Oxford University’s COVID-19 project (Hale et al., 2020), a neural network model was developed to predict the course of the pandemic; this model was then used as a surrogate in evolving another neural network to prescribe NPIs. It discovered different strategies at different phases of the pandemic, such as focusing on schools and workplaces early on, alternating policies over time, and focusing on public information and masking later. This application demonstrated the power of ESP in discovering good tradeoffs for complex real-world decision-making tasks.
@@
Encouraged by this application, the XPRIZE Pandemic Response Challenge (XPRIZE, 2023) was developed in 2020-2021 for NPI optimization using the same predictor+prescriptor approach. The competitors used several different machine learning or other techniques, and drew significant attention to AI-supported decision-making. The results were effective and in some cases used to inform decision-makers (e.g. on the 2021 surge in Valencia, Spain, and on the fall 2021 school openings in Iceland). An important conclusion was that it was possible to convene a group of experts—data scientists, epidemiologists, and public health officials—to build a useful set of tools to advise decision-makers on how to cope with and plan around a health disruption to society.
@@
This experience with the XPRIZE Pandemic Response Challenge then led to Project Resilience (ITU, 2023). The XPRIZE competition serves as a blueprint for a collaborative effort to tackle global problems. The goal of Project Resilience is to build a public AI utility where a global community of innovators and thought leaders can enhance and utilize a collection of data and AI approaches to help with better preparedness, intervention, and response to environmental, health, information, social equity, and similarly scoped problems. The project is hosted by the ITU agency of the United Nations, under its Global Initiative for AI and Data Commons, and contributes to its general efforts toward meeting the Sustainable Development Goals (SDGs).
@@
  • Build an application programming interface (API) for third parties to submit models.
@@
  • Build a portal to visualize the assessment of predictors and prescriptors to include generations of key performance indicators (KPIs) and comparison across models;
@@
3. Land-Use Optimization Task
@@
3.1. Data
The data for carbon emissions (Emissions resulting from Land-Use Change, ELUC) originate from the Bookkeeping of Land-Use Emissions model (BLUE; Hansis et al., 2015), which attributes carbon fluxes to land-use activities. BLUE is one of three bookkeeping models that are used to provide ELUC estimates for the annually published Global Carbon Budget (GCB; Friedlingstein et al., 2023). Out of these three models, BLUE is the only one that provides spatially explicit ELUC estimates. In this paper, BLUE is used to estimate committed emissions due to land-use changes, which means that all the emissions or removals caused by a land-use change event are attributed to the year of the event. The emission (or removal) estimates thus relate to the long-term impact of a land-use change event, i.e. up to multiple decades (Hansis et al., 2015; Houghton et al., 1983). While in principle BLUE can be used as the surrogate model for ESP, in practice its calculations are too expensive to carry out on demand during the search for good policies. Therefore, we performed a number of BLUE simulations covering a comprehensive set of land-use changes for 1850-2022, resulting in a dataset that could be used to train an efficient surrogate model.
The Land-Use Change (LUC) data is provided by the Land-Use Harmonization project (LUH2; Hurtt et al., 2020). LUH2 provides data on fractional land-use patterns, underlying land-use transitions, and key agricultural management information, annually for the time period 850-2100 at 0.25$\times$0.25 degree resolution. LUH2 provides separate layers for states, transitions, and management. Version LUH2-GCB2022 from the Global Carbon Budget 2022 (Friedlingstein et al., 2022) was the basis for this paper. It is based on the HYDE3.3 historical database Klein Goldewijk et al. (2017); Kees Klein Goldewijk and van Zanden (2017) and uses the 2021 wood harvest data provided by the FAO agency of the United Nations. It consists of the following land-use types:
@@ 3.1. Data
    crop: All types pooled into one category
@@
    range: Grazed natural land (i.e., grazing on grassland, savannah, etc.)
BLUE calculates yearly ELUC estimates based on the land-use changes provided by LUH2. Urban areas are assumed to have the same carbon content as croplands. All crop types are currently pooled into a single category. BLUE considers that the degradation of primary to secondary land often lowers the carbon content in vegetation and soils, and it considers pastures and the degradation of natural grasslands and savannahs by grazing. A complete overview of the model can be found in Hansis et al. (2015).
@@
3.2. Decision-making Problem
The modeling approach aims to understand the domain in two ways: (1) In a particular situation, what are the outcomes of the decision maker’s actions? (2) What are the actions that result in the best outcomes, i.e. the lowest carbon emission (or highest removals) with the smallest change in land use? The data is thus organized into context, action, and outcome variables.
Context describes the problem the decision maker is facing, i.e. a particular grid cell, a point in time when the decision has to be made, and the current land use at that time. More specifically it consists of:
@@ 3.2. Decision-making Problem
  • Land use, representing the percentage of land in the cell used by each type above. For simplicity, all crop types were pooled into a single type. In addition, a ‘nonland’ type represents the percentage of the cell that is not land (e.g. typically sea, lake, etc.). The percentages sum up to 100% in each grid cell.
@@
First, decision-makers cannot affect primary land. The idea is that it is always better to preserve primary vegetation, such as pristine rainforest; destroying it is not an option given to the system. Technically, it is not possible to re-plant primary vegetation. Once destroyed, it is destroyed forever. If re-planted, it would become secondary vegetation, for which BLUE assumes lower carbon content than for primary vegetation Hansis et al. (2015).
@@
Second, decision-makers cannot affect urban areas. The needs of urban areas are dictated by other imperatives, and optimized by other decision makers. Therefore, the ESP system cannot recommend that a city should be destroyed, or expanded.
@@
Outcomes consist of two conflicting variables. The primary variable is ELUC, i.e. emissions from land-use change. It consists of all CO2 emissions attributed to the land-use change, in metric tons of carbon per hectare (tC/ha), obtained from the BLUE simulation. Note that all follow-up CO2 emissions (or removals) are attributed to the year of the land-use change, thus neglecting temporal dynamics in CO2 fluxes. A positive number means carbon is emitted to the atmosphere and a negative number means carbon is taken up by vegetation. The secondary variable is the cost of the land-use change, represented by the percentage of land that was changed. This variable is calculated directly from the actions.
@@
There is a tradeoff between these two objectives: It is easy to reduce emissions by changing most of the land, but that would come at a huge cost. Therefore, decision-makers have to minimize ELUC while minimizing land-use change at the same time. Consequently, the result is not a single recommendation, but a Pareto front where each point represents the best implementation of each tradeoff given a balance between the two outcomes.
Model Time (s) EU SA US Global
LinReg (EU) 0.085 0.033 0.171 0.169 0.206
LinReg (SA) 0.570 0.136 0.153 0.062 0.111
LinReg (US) 0.343 0.140 0.146 0.035 0.073
LinReg (Global) 6.872 0.138 0.150 0.036 0.074
RF (EU) 16.493 $0.052$ $0.208^{{\dagger}}$ $0.153^{*{\dagger}}$ $0.220^{{\dagger}}$
RF (SA) 206.451 $0.133^{{\dagger}}$ $\textbf{0.063}^{*{\dagger}}$ $0.076^{{\dagger}}$ $0.127^{{\dagger}}$
RF (US) 101.790 $0.142$ $0.184^{{\dagger}}$ $\textbf{0.019}^{*{\dagger}}$ $0.092^{{\dagger}}$
RF (Global) 422.284 $0.035^{*{\dagger}}$ $0.067^{*{\dagger}}$ $0.020^{*{\dagger}}$ $\textbf{0.040}^{*{\dagger}}$
NeuralNet (EU) 10.650 $\textbf{0.024}^{*{\dagger}}$ $0.266$ $0.282$ $0.333$
NeuralNet (SA) 96.430 $0.252$ $0.099^{*}$ $0.551$ $0.412$
NeuralNet (US) 71.976 $0.127^{*{\dagger}}$ $0.232$ $0.024^{*}$ $0.147$
NeuralNet (Global) 1047.710 $0.045^{*}$ $0.109^{*}$ $0.025^{*}$ $0.048^{*}$
Bold: the best model in each region; *: 99% confidence the model outperforms LinReg; †: 99% confidence that RF outperforms NeuralNet, or vice-versa
Table 1. Average training time and mean absolute error in tC/ha of the trained models of each type trained on each region and evaluated on that region as well as all other regions. Statistical significance was estimated in each case with the unpaired $t$-test. The RF models are more accurate than NeuralNets (with the exception of EU), which are more accurate than LinReg, and overall the specialized models are more accurate than the Global models. However, RF models do not extrapolate well (as seen in Figure 2), which is necessary for evaluating prescriptor candidates. In contrast, NeuralNets do, and they are also otherwise reasonably accurate. Therefore, the Global NeuralNet model was used to evolve the prescriptors.
4. Models
@@
The system consists of the predictor, trained with supervised learning on the historical data, and the prescriptor, evolved against the predictor.
@@
Prediction: Given the context and actions that were performed, the predictive model estimates the outcomes. In this case, since the cost outcome (i.e. the percentage of changed land) can be calculated directly, only the ELUC is predicted by the model. That is, given the land use of a specific location, and the land-use changes that were made during a specific year, the model predicts the long-term CO2 emissions directly caused by these land-use changes.
@@
Any predictive model can be used in this task, including linear regression, random forest, or a neural network; they will each be evaluated below. As usual in machine learning, each model is fit to the existing historical data and evaluated with left-out data.
@@
Prescription: Given context, the prescriptive model suggests actions that optimize the outcomes. The model has to do this for all possible contexts, and therefore it represents an entire strategy for optimal land use. The strategy can be implemented in various ways, like sets of rules or neural networks. The approach in this paper is based on neural networks.
@@ 5. Experiments
This section describes the main results of the project so far, including different ways to create good predictors and the evolution of effective prescriptors based on them. The behavior of the discovered prescriptors is outlined, and an interactive demo of the entire land-use recommendation system is described.
5.1. Prediction